OpenStudy (anonymous):

x2 + 5x + 6 = 0 A. {2, 3} B. {1, –6} C. {–1, 6} D. {–2, –3}

jagr2713 (jagr2713):

Hey !!! Which two numbers add up to 5 and multiply to 6? 2 and?

OpenStudy (anonymous):

3

OpenStudy (phebe):

(x+2) (x+3) = 0

jagr2713 (jagr2713):

Good so we rewrite the expression as (x+?)(x+?)

jagr2713 (jagr2713):

Well seems like @Phebe already said it so : When will (x+2)(x+3) equal zero?

OpenStudy (anonymous):

with -2 and -3

jagr2713 (jagr2713):

its when x+2=0 and?

OpenStudy (anonymous):

option D

OpenStudy (anonymous):

so d!

jagr2713 (jagr2713):

@danial.sheikh It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

jagr2713 (jagr2713):

Correct b.c its when x+2=0 and x+3=0 And we solve so we subtract 2 and 3 on both sides and get -2,-3

OpenStudy (anonymous):

can you help with more?

OpenStudy (anonymous):

What value should be added to the expression to create a perfect square? x2 – 10x A. 25 B. –100 C. 5 D. 100

jagr2713 (jagr2713):

One question per post